# Slow magnetic quantum oscillations in the $c$-axis magnetoresistance of UTe2

### v1

@@ I Introduction
Key to unraveling the nature of novel quantum phases in unconventional metals is a detailed knowledge about their electronic band structure [1]. Since the discovery of unconventional superconductivity in UTe2 in 2019 [2, 3], evidence for a spin-triplet ground state in this material has been reported from various experiments. For example: Enhanced upper critical fields beyond the Pauli-paramagnetic limit were observed [4]; an anisotropic spin susceptibility was probed by nuclear magnetic resonance spectroscopy [5, 6, 7, 8]; high magnetic field re-entrant superconductivity was detected in pulsed-field experiments surviving up to 70 T for specific directions [9, 10, 11, 12, 13]. Yet, the overall ground state of UTe2 remains enigmatic and under hot debate [4, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], notably regarding the potential of a topological chiral superconducting ground state [9, 12, 11]. Moreover, novel mechanisms may play a leading role in its field-enhanced superconducting state with distinct order parameters between the low field-low pressure phase and the high field-high pressure phase [9, 27, 12, 28]. Magnetic field and its orientation with respect to the crystal lattice plays an important role in the behavior of UTe2. Field-induced changes of the Fermi surface (FS) are indeed expected for different field orientations: A metamagnetic transition was observed for a field above 35 T aligned along the $b$ axis [29, 30, 31, 32, 2]; changes in thermo-electric transport indicate a potential Lifshitz transition at 6 T for $H\parallel a$ [33].
@@
The electronic band structure of UTe2 has been intensely investigated by multiple methods such as photoemission spectroscopy [34, 35, 36, 23] and x-ray spectroscopy [37, 38, 39], providing first insights on the FS and the influence of $5f$ electrons in UTe2. All experiments point to an intermediate valence state close to 2.5, even though the degree of localization and the exact valence state of the uranium in UTe2 are still under debate. Besides two cylindrical electron- and hole-like quasi-two-dimensional (quasi-2D) FS sections, additional 3D FS pockets were reported [34, 35, 40]. Magnetic quantum oscillations (MQOs), so-called de Haas-van Alphen (dHvA) oscillations, detected by torque magnetometry [41] and Shubnikov-de Haas (SdH) oscillations, detected by proximity-/tunnel-diode-oscillator (P/TDO) measurements [42, 40] and magnetoresistance (MR) [43], confirmed the presence of warped 2D FS cylinders. Moreover, additional low-frequency MQOs were detected by the PDO and TDO method that do not match the large quasi-2D FSs [40, 44]. These low-frequency MQOs might be associated with small FS pockets. The presence or absence of such a small FS pocket, with heavy or light band masses, is essential for understanding the superconducting order parameter in UTe2 [25, 45]. The experimental evidence for such small FSs, however, is being discussed controversially: While one experiment implies a 3D pocket from the detection of slow MQOs that were observed for any orientation of the external field [40], others found signatures of such low frequencies only for particular directions [44]. In the latter case, the MQOs have been associated with quantum interference between sections of two distinct quasi-2D FS cylinders [44]. Quantum interference arises from coherent charge-carrier pathways between different magnetic-breakdown (MB) junctions, effectively enclosing flux in an interferometer-like configuration [46, 47, 48]. MB is a well-known effect in magnetic field based on a tunnel transition of conduction electrons of a metal between trajectories belonging to different electronic bands [49, 1]. According to Ref. [49], the observation of quantum interference requires a FS that (i) realizes an interferometer setup, i.e., at least two semiclassical paths on which particles travel from one MB junction to the next and which enclose an area in momentum space; (ii) MB junctions between the semiclassical paths with intermediate MB probabilities $p$ leading to a partial MB with $p(1-p)\neq 0$. One of the key signatures of quantum interference oscillations (QIOs) is their stability against increased temperature [1]. According to the theory, the effective mass of QIOs results from the difference of the two involved parts of the FS. Hence, it can be significantly smaller than compared to the fundamental renormalized band masses, reducing the damping of the oscillation amplitude due to temperature. Specifically, this feature may enable a detection of the MQOs at higher temperatures in the case of the heavy fermion UTe2 [42].
@@
Nevertheless, alternative origins for the observed MQOs may still be possible for quasi-2D FSs: For example, forbidden orbits from oscillations in the chemical potential [50], quasiparticle-lifetime oscillations (QPLOs) [51, 48], and slow oscillations from weakly warped 2D FSs [52, 53] may result in rather low-effective-mass frequencies similar to what is reported for UTe2. Moreover, the presence of additional 3D FS pockets may enable more complex scenarios for possible trajectories via the MB mechanism observable in MQO studies.
@@
To date, the low-frequency quantum oscillations have only been observed in the electrical transport of UTe2 and not in dHvA measurements  [41, 43]. Notably, the previous reports on the observation of slow oscillations from small FS sheets rely on results obtained with the P/TDO technique, which mainly is sensitive to the conductivity on the surface of the sample. It is, therefore, essential to confirm the reported MQOs with a bulk method.
@@
Here, we investigate the angle-resolved magnetoresistance and Hall effect in bulk crystalline samples with current along the $c$ axis in fields up to 606060\,60T. We investigate the temperature and angle dependence of low-frequency MQOs of single-crystalline UTe2 in steady and pulsed magnetic fields up to 24 and 60 T, respectively. We find two distinct frequencies, $F_{1}\approx 90\,$T and $F_{2}\approx 205\,$T, that emerge consecutively. These oscillations show up for field aligned within a narrow angular region around the crystallographic $a$ direction. $F_{2}$ matches previously reported frequency values obtained from P/TDO measurements associating the oscillations with quantum interference. Our results imply yet unreported Lifshitz transitions that affect the Fermi-surface topology. We discuss the origin of these oscillations in terms of magnetic breakdown (MB), quantum interference, and other potential mechanisms.
@@ II Methods
Crystal growth: UTe2 single crystals with $T_{\mathrm{c}}$ $=2.1\,$K were prepared by the molten-salt-flux (MSF) growth technique at Tohoku University, Oarai (Japan) and CEA, Grenoble (France). Sample S1 is a bulk sample, polished down to a thickness of about 100100100\,100µm, which was contacted using gold wires and spot welding. The contacts showed less than $1\,\Omega$ resistances. Sample S2 is a focused-ion-beam (FIB) microfabricated Hall-bar structure (with a $T_{\mathrm{c}}$ of 2.1 K) cut from a single crystal similar to that reported in Ref. [12]. In order to prevent degradation due to exposure to air, the structure, including the contacts, was encased in insulating epoxy. The cross section of S2 is $(w\times t)=(2.0\times 4.0)$ µm2 with a length of $l=24\,$µm between the voltage contacts.
@@
DFT calculations: We reproduced previous density functional theory (DFT) calculations [15] utilizing the full-potential local orbit minimum basis code (FPLO, version 22.00-26) [54]. We used the general gradient approximation (GGA) [55] and applied an additional, repulsive Coulomb potential on the $5f$ electrons of the uranium atoms a few eV (GGA+U). This potential moves the otherwise itinerant $5f$ electrons into localized states. We conducted all calculations fully relativistic to take the impact of spin-orbit coupling into account. We used a $k$-mesh with $20\times 20\times 20$ points to calculate the density of states and band structure and a finer grid with 40 (80 for $U=2\,$eV) points per dimension for calculating the FS.
@@ III Results and Discussion
Figure 1: (a) $c$-axis MR ($I\parallel c$) of the bulk sample S1 recorded at 0.5 K for $H\parallel a$ with $I=1\,$mA. Inset: Residual resistance after subtraction of a fit with polynomial of $5^{\mathrm{th}}$ degree (red line in main panel) to the data between 10 and 24 T. (b) Residual resistance after background subtraction (polynomial of $5^{\mathrm{th}}$ degree) for $H\parallel a$ and $T=0.5$ and 1.41.41.4\,1.4K plotted against inverse magnetic field. Blue and red dashed lines indicate the $1/(\mu_{0}H)$ oscillation periods for the respective frequencies $F_{1}$ and $F_{2}$.
We investigated two different samples of similar high quality labeled S1 and S2 (for sample details see experimental section above). We measured the MR of S1 at various temperatures between 0.6 and 333\,3K, with $I\parallel c$ and field oriented along $a$. The overall normal-state MR increases monotonically with increasing field at least at low temperature, see Fig. 1(a). Slow SdH oscillations, discernible already with bare eye, emerge above 10 T [see also the background-subtracted residual-resistance curves in the inset of Fig. 1(a) and in Fig. 1(b)]. Apparently, there are two distinct oscillation periods, $1/F_{1}$ and $1/F_{2}$ [red and blue dashed lines in Fig. 1(b)], indicative of more than one oscillation frequency, but emerging consecutively during the field sweep below and above 16 T, respectively. In general, MQOs are expected to be periodic in $1/H$. However, we find that beyond 16 T, $1/F_{1}$ does not match with the oscillation maxima/minima any longer and the smaller amplitude oscillations with $1/F_{2}$ emerge.
In Figs. 2(a) and 2(b), we present further MR data recorded for the two samples, S1 and S2, at various temperatures ranging from 0.5 to 2.5 K. For both samples, a negative MR with a hump-like feature at around $6-7\,$T appears in the normal state upon increasing the temperature. This hump has already been observed in previous magnetotransport studies that connected it to a feature observed simultaneously in Hall as well as in the Seebeck effect [33]. It was ascribed to a Lifshitz transition, which suggests a significant change in the electronic band structure. In the fast-Fourier-transformation (FFT) spectra for the low- and high-field range [Fig. 2(c)], peaks at $F_{1}\approx 85$ T and at $F_{2}\approx 200$ T show up. Note: As the two oscillations exhibit different amplitudes and are best discernible within separate field ranges below or above 16 T, we chose two distinct field windows, namely $10-14\,$T and $17-24\,$T, for the analyses of the temperature damping of $F_{1}$ and $F_{2}$, respectively. Moreover, we describe the temperature dependence of the FFT amplitudes according to the Lifshitz-Kosevich (LK) damping term:
@@
with $\alpha=-14.69\,$T/K and $m^{*}$ being the effective electron mass. Both frequencies are damped almost equally by temperature, indicating rather low effective-mass values of approximately $1.5(2)m_{\mathrm{e}}$, where $m_{\mathrm{e}}$ is the free-electron mass. The fits of the mass plots are presented in Fig. 2(e), left panel. We, furthermore, conducted measurements in a 15 T magnet system and confirmed the effective mass value for S1 with $m_{1}^{*}\approx 1.5(1)m_{\mathrm{e}}$ for $F_{1}$ [see Fig. 2(e), right panel].
@@
Figure 2: (a), (b) $c$-axis resistance of S1 and S2, respectively, recorded in two different setups at various temperatures for field aligned parallel to the $a$ axis. (c) FFT spectra of the residual resistance of S1 obtained from the data shown in (a) for two different field ranges, $10-14\,$T and $17-24\,$T, respectively. (d) Residual background-subtracted MR of S2 from data in (b) plotted against magnetic field. (e) Effective-mass plots, Left: FFT amplitudes for S1 versus temperature for $F_{1}$ and $F_{2}$ for two field windows $10-14\,$T and $17-24\,$T, respectively. Right: Amplitude at 11.8 T for S1. (f) Effective-mass plot for the mean amplitude values between the maximum and minimum around 14.2 T according to Eq. (1) for S2.
@@
A similar MQO as $F_{1}$ is observable in the transport signal up to 16 T of the FIB microfabricated sample S2 [Fig. 2(d)]. The determined effective-mass value is $1.5(1)m_{\mathrm{e}}$ [cf. mass plot in Fig. 2(f)], so basically the same as obtained for S1. Even though the microstructured sample exhibits lower critical fields for $H\parallel a$, the oscillations look very similar with approximately the same low frequency of about $F_{1}=90\,$T. Interestingly, for S2, we observe an additional higher-order variation in the residual resistance, visible by two local maxima in the oscillation near 12 T [Fig. 2(d)]. The split maximum seems not to be present in the data for S1. Assuming a periodicity according to the frequency $F_{2}$, the split maximum at 12 T cannot be associated with that frequency. As we already described above for sample S1, the periodicity of $F_{1}$ seems to change in the vicinity of 16 T, just before it is suppressed completely. Hence, the variation in the periodicity of $F_{1}$ for sample S2 may have a similar origin.
@@
Figure 3: (a) MR of S2 recorded at 1.5 K in a pulsed magnet up to 60 T with $H\parallel a$. A slight heating, likely caused by eddy currents, is responsible for the deviation between the up and down sweep. (b) Residual resistance after background subtraction of a cubic polynomial. (c) FFT spectra for a field window from 20 to 59 T. Note: The peak at 50 T strongly depends on the choice of the polynomial degree and is therefore not related to MQOs.
@@
In an attempt to trace the oscillations towards higher fields, we studied S2 down to 1.5 K in a pulsed-field setup up to 60 T, see Fig. 3. Owing to the enhanced contact resistances, the sample received a significant heat load from the application of a current of 100100100\,100µA required for a reasonable signal-to-noise ratio. The comparison to the steady-field data indicates a sample temperature of about 2.5 K. In the pulsed-field MR data in Fig. 3(a), we see a deviation between the up and down sweep at low fields that originates from additional heating during the pulse, potentially related to eddy currents. Moreover, the high-resistance contacts forced us to apply a relatively low excitation frequency of 11 kHz. With an approximate pulse rise time of 30 ms and a pulse decline within 100 ms, this further limited the resolution during the experiment. Nevertheless, we were able to resolve MQOs as can be seen from Fig. 3(b). For $H\parallel a$, the FFT spectrum exhibits a clear peak at around $F=210(5)\,$T, which agrees within error bars with the value of $F_{2}$ determined for S1 in steady fields up to 24 T.
@@
Figure 4: (a),(b) Residual MR (background subtraction using a polynomial of $5^{\mathrm{th}}$ degree) recorded for S1 at 1.41.41.4\,1.4K with $I=2\,$mA for various tilt angles within the $a$-$b$ and $a$-$c$ rotation plane, respectively. Note: Curves in (a)-(d) are shifted by constant offsets for better visibility. (c),(d) Residual MR (background subtraction using a polynomial of $5^{\mathrm{th}}$ degree) for S2 measured at various tilt angles at 0.5 and 0.03 K, respectively, with $I=0.02\,$mA, recorded during two separate experiments in a 16 T and an 18 T magnet system. (e) FFT spectra for S1 and S2 and $H\parallel a$. (f) Frequency values determined from the periodicity of the oscillations as indicated by circles and squares in (a)-(d) for sample S1 (black circles and red squares) and S2 (blue triangles). The blue diamond marks the data point obtained in 60 T for $H\parallel a$.
@@
In the following, we will turn to the angle dependence of the observed MQOs. We measured MR of S2 up to 60 T for $H\parallel b$, $H\parallel c$, and $40^{\circ}$ tilted away from $b$ towards $c$. We could not resolve any MQOs, at least within the noise level that was similar to the one in Fig. 3. We were able to trace oscillations within the $a$-$b$ (azimuthal angle $\phi$) and the $a$-$c$ (polar angle $\theta$) plane for both samples in steady fields in separate setups and laboratories. In Figs. 4(a)-4(d), we present selected residual MR curves obtained at various tilt angles. In order to trace the frequencies $F_{1}$ and $F_{2}$ for different angles, we extracted the periodicity between maxima and minima (marked by black circles and squares) and plot them for both samples in Fig. 4(f). As can be seen in Fig. 4(e), an analysis based on a FFT yields rather broad peaks due to the low number of maxima and minima within the given field ranges. Moreover, FFTs are very sensitive to the background subtraction and the degree of the chosen polynomial, leading to a smearing out of the FFT peaks in the frequency range of 0 to approximately 50 T. Hence, we determined the periodicity $1/F$ between the observed maxima, as indicated by the black circles in Figs. 4(a)-4(d)
@@
While a frequency of about 90 T has not been reported to date, $F_{2}$ matches the 200 T frequency previously observed in the surface conduction for $H\parallel a$ [42, 44]. There, these MQOs have been associated with Stark oscillations that are caused by quantum interference as introduced above [47].
Figure 5: (Black curve in both panels) Same data as shown in Fig. 1(b). (Blue curve) Simulation according to the LK formula Eq.(1) with $F=85\,$T and $m^{*}=1.5\,m_{\mathrm{e}}$. (Red curve) Simulation according to the LK formula with $F=205\,$T and $m^{*}=1.5\,m_{\mathrm{e}}$. (Magenta curve) Superposition of both frequencies with parameters from both simulations above, plus MB damping factors $R_{1,\mathrm{MB}}$ and $R_{2,\mathrm{MB}}$, with $B_{\mathrm{MB}}=6\,$T.
@@
$\Delta\rho=B^{1/2}\sum_{i=1,2}A_{i}R_{i,\mathrm{D}}R_{i,\mathrm{MB}}R_{\mathrm{i,S}}\cos{\left(2\pi\left(\frac{F_{i}}{B}+\gamma_{i}\right)\right)},$ (2)

### v2

@@ I Introduction
Key to unraveling the nature of novel quantum phases in unconventional metals is a detailed knowledge about their electronic band structure [1]. Since the discovery of unconventional superconductivity in UTe2 in 2019 [2, 3], evidence for a spin-triplet ground state in this material has been reported from various experiments. A few examples of these are as follows: Enhanced upper critical fields beyond the Pauli-paramagnetic limit were observed [4]; an anisotropic spin susceptibility was probed by nuclear magnetic resonance spectroscopy [5, 6, 7, 8]; and high magnetic-field re-entrant superconductivity was detected in pulsed-field experiments surviving up to 70 T for specific directions [9, 10, 11, 12, 13]. Yet, the overall ground state of UTe2 remains enigmatic and under intense debate [4, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], notably regarding the potential of a topological chiral superconducting ground state [9, 12, 11]. Moreover, novel mechanisms may play a leading role in its field-enhanced superconducting state with distinct order parameters between the low field-low pressure phase and the high field-high pressure phase [9, 27, 12, 28]. Magnetic field and its orientation with respect to the crystal lattice play an important role in the behavior of UTe2. Field-induced changes of the Fermi surface (FS) are indeed expected for different field orientations: A metamagnetic transition was observed for a field above 35 T aligned along the $b$ axis [29, 30, 31, 32, 2]; changes in thermo-electric transport indicated a potential Lifshitz transition at 6 T for $H\parallel a$ [33].
The electronic band structure of UTe2 has been intensely investigated by multiple methods such as photoemission spectroscopy [34, 35, 36, 23] and x-ray spectroscopy [37, 38, 39], providing first insights on the FS and the influence of $5f$ electrons in UTe2. All experiments point to an intermediate valence state close to 2.5, even though the degree of localization and the exact valence state of the uranium in UTe2 are still under debate. Besides two cylindrical electron- and hole-like quasi-two-dimensional (quasi-2D) FS sections, additional three-dimensional (3D) FS pockets were reported [34, 35, 40].
@@
Magnetic quantum oscillations (MQOs), so-called de Haas–van Alphen (dHvA) oscillations, detected by torque magnetometry [41] and Shubnikov–de Haas (SdH) oscillations, detected by proximity-/tunnel-diode-oscillator (P/TDO) measurements [42, 40] and magnetoresistance (MR) [43], confirmed the presence of warped 2D FS cylinders. Moreover, additional low-frequency MQOs were detected by the PDO and TDO method that did not match the large quasi-2D FSs [40, 44]. These low-frequency MQOs might be associated with small FS pockets. The presence or absence of such a small FS pocket, with heavy or light band masses, is essential for understanding the superconducting order parameter in UTe2 [25, 45]. The experimental evidence for such small FSs, however, is being discussed controversially: While one experiment implied a 3D pocket from the detection of slow MQOs that were observed for any orientation of the external field [40], others found signatures of such low frequencies only for particular directions [44]. In the latter case, mainly due to the corresponding low effective masses, the slow MQOs have been associated with quantum interference between sections of two distinct quasi-2D FS cylinders [44].
@@
Nevertheless, alternative origins for the observed MQOs may still be possible for quasi-2D FSs: One of the most prominent effects that can lead to the emergence of additional frequencies is magnetic breakdown between trajectories belonging to different electronic bands [1]. Also, more exotic effects, such as oscillations in the chemical potential [46], quasiparticle-lifetime oscillations [47, 48], or slow oscillations from weakly warped 2D FSs [49, 50], may account for low-effective-mass frequencies similar to what is reported for UTe2. Moreover, the presence of additional 3D FS pockets may enable more complex scenarios for possible trajectories via the MB mechanism observable in MQO studies.
@@
To date, the low-frequency quantum oscillations have only been observed in the electrical transport of UTe2 and not in dHvA measurements  [41, 43]. Notably, the previous reports on the observation of slow oscillations from small FS sheets relied on results obtained with the P/TDO technique, which mainly were sensitive to the conductivity on the surface of the sample. It is, therefore, essential to confirm the reported MQOs with a bulk method.
@@
Here, we investigate the angle-resolved magnetoresistance and Hall effect in bulk crystalline samples with current along the $c$ axis in fields up to 606060\,60T. We investigate the temperature and angle dependence of low-frequency MQOs of single-crystalline UTe2 in steady and pulsed magnetic fields up to 24 and 60 T, respectively. We find two distinct frequencies, $F_{1}\approx 90\,$T and $F_{2}\approx 205\,$T, that emerge consecutively. These oscillations show up for a field aligned within a narrow angular region around the crystallographic $a$ direction. $F_{2}$ matches the previously reported frequency values obtained from P/TDO measurements associating the oscillations with quantum interference. Our results imply as-yet unreported Lifshitz transitions that affect the Fermi-surface topology. We discuss the origin of these oscillations in terms of magnetic breakdown (MB), quantum interference, and other potential mechanisms.
@@ II Methods
Crystal growth: UTe2 single crystals were prepared by the molten-salt-flux (MSF) growth technique at Tohoku University, Oarai (Japan) and CEA, Grenoble (France), respectively. Sample S1 is a bulk sample with $T_{\mathrm{c}}$ $=2.1\,$K, polished down to a thickness of about 100100100\,100µm, which was contacted using gold wires and spot welding, exhibiting a residual resistance ratio (RRR) of about 24. The contacts showed less than $1\,\Omega$ resistances. Sample S2 is a focused-ion-beam (FIB) microfabricated Hall-bar structure (with a $T_{\mathrm{c}}$ of 1.9 K and RRR $\approx 15$) cut from a single crystal with $T_{\mathrm{c}}$ $=2\,$K, determined by specific-heat measurements. To prevent degradation due to exposure to air, the structure, including the contacts, was encased in an insulating epoxy. The encapsulation may introduce additional pressure to the structure that could be the reason for the slightly reduced $T_{\mathrm{c}}$. The cross section of S2 was $(w\times t)=(2.0\times 4.0)$ µm2 with a length of $l=24\,$µm between the voltage contacts. A second sample, S3, was cut from the same crystal with $I\parallel a$ that exhibited a $T_{\mathrm{c}}$ of about 2 K and a RRR of 105 in resistance measurements.
@@
DFT calculations: We reproduced previous density functional theory (DFT) calculations [15] utilizing the full-potential local orbit minimum basis code (FPLO, version 22.00-26) [51]. We used the general gradient approximation (GGA) [52] and applied an additional, repulsive Coulomb potential on the 5f electrons of the uranium atoms a few eV. This potential moved the otherwise itinerant $5f$ electrons into localized states. We conducted all calculations as fully relativistic to take the impact of spin-orbit coupling into account. We used a $k$-mesh with $20\times 20\times 20$ points to calculate the density of states and band structure and a finer grid with 40 (80 for $U=2\,$eV) points per dimension for calculating the FS.
@@ III Results and Discussion
Figure 1: (a) $c$-axis MR ($I\parallel c$) of the bulk sample S1 recorded at 0.5 K for $H\parallel a$ with $I=1\,$mA. Inset: Residual MR after subtraction of a fit with polynomial of fifth degree (red line in main panel) to the data between 10 and 24 T. (b) Residual MR after background subtraction (polynomial of fifth degree) for $T=0.5$ and 1.41.41.4\,1.4K, respectively, plotted against inverse magnetic field $H\parallel a$. Blue and red dashed lines indicate the $1/(\mu_{0}H)$ oscillation periods for the respective frequencies $F_{1}$ and $F_{2}$.
We investigated two different samples of a similar high quality, labeled S1 and S2 (for sample details see the experimental section above). We measured the MR of S1 at various temperatures between 0.6 and 333\,3K, with $I\parallel c$ and the field oriented along $a$. The overall normal-state MR increased monotonically with increasing field, at least at low temperature, see Fig. 1(a). Slow SdH oscillations, discernible already with the bare eye, emerged above 10 T [see also the background-subtracted residual-MR curves in the inset of Fig. 1(a) and in Fig. 1(b)]. Apparently, there are two distinct oscillation periods, $1/F_{1}$ and $1/F_{2}$ [red and blue dashed lines in Fig. 1(b)], indicative of more than one oscillation frequency, but emerging consecutively during the field sweep below and above 16 T, respectively. In general, MQOs are expected to be periodic in $1/H$. However, we find that beyond 16 T, $1/F_{1}$ does not match with the oscillation maxima/minima any longer and the smaller amplitude oscillations with $1/F_{2}$ emerge.
In Figs. 2(a) and 2(b), we present further MR data recorded for the two samples, S1 and S2, at various temperatures ranging from 0.5 to 2.5 K. For both samples, a negative MR with a hump-like feature at around 6-7 T appears in the normal state upon increasing the temperature. This hump was already observed in previous magnetotransport studies that connected it to a feature observed simultaneously in the Hall as well as in the Seebeck effects [33]. It was ascribed to a Lifshitz transition, which suggests a significant change in the electronic band structure. In the fast-Fourier-transformation (FFT) spectra for the low- and high-field range [Fig. 2(c)], peaks at $F_{1}\approx 85$ T and at $F_{2}\approx 200$ T show up. Note that as the two oscillations exhibit different amplitudes and are best discernible within separate field ranges below or above 16 T, we chose two distinct field windows, namely, 10-14 T and 17-24 T, for the analyses of the temperature damping of $F_{1}$ and $F_{2}$, respectively. Moreover, we describe the temperature dependence of the FFT amplitudes according to the Lifshitz-Kosevich (LK) damping term:
@@
with $\alpha=-14.69\,$T/K and $m^{*}$ being the effective electron mass. Both frequencies are damped almost equally by temperature, indicating low effective-mass values of approximately $1.5(2)m_{\mathrm{e}}$, where $m_{\mathrm{e}}$ is the free-electron mass. The fits of the mass plots are presented in Fig. 2(e), left panel. We, furthermore, conducted measurements in a 15 T magnet system and confirmed the effective mass value for S1 with $m_{1}^{*}\approx 1.5(1)m_{\mathrm{e}}$ for $F_{1}$ [see Fig. 2(e), right panel].
@@
Figure 2: (a), (b) $c$-axis resistance of S1 and S2, respectively, recorded in two different setups at various temperatures for field aligned parallel to the $a$ axis. (c) FFT spectra of the residual MR of S1 obtained from the data shown in (a) for two different field ranges, 10-14 T and 17-24 T, respectively. (d) Residual background-subtracted MR (fifth degree polynomial) of S2 from data in (b) plotted against magnetic field. (e) Effective-mass plots, Left: FFT amplitudes for S1 versus temperature for $F_{1}$ and $F_{2}$ for two field windows 10-14 T and 17-24 T, respectively. Right: Amplitude at 11.8 T for S1. (f) Effective-mass plot for the mean amplitude values between the maximum and minimum around 14.2 T according to Eq. (1) for S2.
@@
A similar MQO to $F_{1}$ is observable in the transport signal up to 16 T of the FIB microfabricated sample S2 [Fig. 2(d)]. The determined effective-mass value is $1.5(1)m_{\mathrm{e}}$ [cf. mass plot in Fig. 2(f)], so basically the same as that obtained for S1. Even though the microstructured sample exhibits lower critical fields for $H\parallel a$, the oscillations look very similar with approximately the same low frequency of about $F_{1}=90\,$T. Interestingly, for S2, we observe an additional higher-order variation in the residual MR, visible by two local maxima in the oscillation near 12 T [Fig. 2(d)]. The split maximum seems not to be present in the data for S1. Assuming a periodicity according to the frequency $F_{2}$, the split maximum at 12 T cannot be associated with that frequency. As we already described above for sample S1, the periodicity of $F_{1}$ seems to change in the vicinity of 16 T, just before it is suppressed completely. Hence, the variation in the periodicity of $F_{1}$ for sample S2 may have a similar origin.
@@
Figure 3: (a) MR of S2 recorded at 1.5 K in a pulsed magnet up to 60 T with $H\parallel a$. A slight heating, likely caused by eddy currents, is responsible for the deviation between the up and down sweep. (b) Residual MR after background subtraction of a cubic polynomial. (c) FFT spectra for a field window from 20 to 59 T. Note: The peak at 50 T strongly depends on the choice of the polynomial degree and is therefore not related to MQOs.
@@
In an attempt to trace the oscillations towards higher fields, we study S2 down to 1.5 K in a pulsed-field setup up to 60 T, see Fig. 3. Owing to the enhanced contact resistances, the sample receives a significant heat load from the application of a current of 100100100\,100µA required for a reasonable signal-to-noise ratio. The comparison to the steady-field data indicates a sample temperature of about 2.5 K. In the pulsed-field MR data in Fig. 3(a), we see a deviation between the up and down sweeps at low fields that originates from additional heating during the pulse, potentially related to eddy currents. Moreover, the high-resistance contacts forced us to apply a relatively low excitation frequency of 11 kHz. With an approximate pulse rise time of 30 ms and a pulse decline within 100 ms, this further limited the resolution during the experiment. Nevertheless, we were able to resolve MQOs, as can be seen from Fig. 3(b). For $H\parallel a$, the FFT spectrum exhibits a clear peak at around $F=210(5)\,$T, which agrees within error bars with the value of $F_{2}$ determined for S1 in steady fields up to 24 T.
@@
Figure 4: (a),(b) Residual MR (background subtraction using a polynomial of fifth degree) recorded for S1 at 1.41.41.4\,1.4K with $I=2\,$mA for various tilt angles within the $a$-$b$ and $a$-$c$ rotation plane, respectively. Note: Curves in (a)–(d) are shifted by constant offsets for better visibility. (c),(d) Residual MR (background subtraction using a polynomial of fifth degree) for S2 measured at various tilt angles at 0.5 and 0.03 K, respectively, with $I=0.02\,$mA, recorded during two separate experiments in a 16 T and an 18 T magnet system. (e) FFT spectra for S1 and S2 and $H\parallel a$. (f) Frequency values determined from the periodicity of the oscillations as indicated by circles and squares in (a)–(d) for sample S1 (black circles and red squares) and S2 (blue triangles). The blue diamond marks the data point obtained in 60 T for $H\parallel a$.
@@
In the following, we will turn to the angle dependence of the observed MQOs. We measured MR of S2 up to 60 T for $H\parallel b$, $H\parallel c$, and $40^{\circ}$ tilted away from $b$ towards $c$. We could not resolve any MQOs, at least within the noise level that was similar to the one in Fig. 3. We were able to trace oscillations within the $a$-$b$ (azimuthal angle $\phi$) and the $a$-$c$ (polar angle $\theta$) plane for both samples in steady fields in separate setups and laboratories. In Figs. 4(a) to 4(d), we present selected residual MR curves obtained at various tilt angles. To trace the frequencies $F_{1}$ and $F_{2}$ for different angles, we extract the periodicity between the maxima and minima (marked by black circles and squares) and plot them for both samples in Fig. 4(f). As can be seen in Fig. 4(e), an analysis based on a FFT yields broad peaks due to the low number of maxima and minima within the given field ranges. This, in particular, leads to a broadening of the FFT peaks in the low-frequency range between 0 and approximately 50 T, which is close to the value of $F_{1}$. Moreover, the amplitudes of the peaks in the FFT spectra are very sensitive to the background subtraction and the degree of the chosen polynomial. We also checked lower-degrees polynomials for subtracting the background and made sure that no frequencies were removed or added. Here, we determine the periodicity $1/F$ between the observed maxima, as indicated by the black circles in Figs. 4(a) to 4(d)
@@
While a frequency of about 90 T has not been reported to date, $F_{2}$ matches the 200 T frequency previously observed in the surface conduction for $H\parallel a$ [42, 44]. There, these MQOs were associated with Stark oscillations that were caused by quantum interference between different sections of the FS [53].
Figure 5: (Black curve in both panels) Same data as shown in Fig. 1(b). (Blue curve) Simulation according to the LK formula, Eq.(1), with $F=85\,$T and $m^{*}=1.5\,m_{\mathrm{e}}$. (Red curve) Simulation according to the LK formula with $F=205\,$T and $m^{*}=1.5\,m_{\mathrm{e}}$. (Magenta curve) Superposition of both frequencies with parameters from both simulations above, plus MB damping factors $R_{1,\mathrm{MB}}$ and $R_{2,\mathrm{MB}}$, with $B_{\mathrm{MB}}=6\,$T.
@@
$\Delta\rho=B^{1/2}\sum_{i=1,2}A_{i}R_{i,\mathrm{D}}R_{i,\mathrm{MB}}R_{\mathrm{i,S}}\cos{\left[2\pi\left(\frac{F_{i}}{B}+\gamma_{i}\right)\right]},$ (2)